# MIMIR: Masked Image Modeling for Mutual Information-based Adversarial Robustness

### v2

@@ Appendix G Mutual Information and HSIC
Unfortunately, estimating MI in high-dimensional space is a difficult task since it may involve a precise estimation of the underlying data distribution $P_{(X,Y)}$ or $P_{(X)}$ and $P_{(Y)}$. To address this issue, the deterministic information bottleneck (DIB) [94] uses the recently proposed matrix-based Rényi’s $\alpha$-entropy functional $I_{\alpha}$ [66, 93], which suggests similar quantities to $I(X,Y)$ in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the reproducing kernel Hilbert space (RKHS), but avoids density estimation.
@@
The Hilbert–Schmidt Independence Criterion (HSIC) [34] is also a kernel-based dependence measure and is usually used as a surrogate of MI. Formally, the HSIC is defined as the squared norm of the cross-covariance operator $||C_{XY}||^{2}$:
@@
in which $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, and $H=I-\frac{1}{N}\mathbbm{1}\mathbbm{1}^{T}$ is the centering matrix.

### v3

@@ Appendix F Mutual Information and HSIC
Unfortunately, estimating MI in high-dimensional space is a difficult task since it may involve a precise estimation of the underlying data distribution $P_{(X,Y)}$ or $P_{(X)}$ and $P_{(Y)}$. To address this issue, the deterministic information bottleneck (DIB) (Yu, Yu, and Príncipe 2021) uses the recently proposed matrix-based Rényi’s $\alpha$-entropy functional $I_{\alpha}$ (Sanchez Giraldo, Rao, and Principe 2015; Yu et al. 2019), which suggests similar quantities to $I(X,Y)$ in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the reproducing kernel Hilbert space (RKHS), but avoids density estimation.
@@
The Hilbert–Schmidt Independence Criterion (HSIC) (Gretton et al. 2005) is also a kernel-based dependence measure and is usually used as a surrogate of MI. Formally, the HSIC is defined as the squared norm of the cross-covariance operator $||C_{XY}||^{2}$:
@@
in which $(K_{x})_{i,j}=\kappa_{x}(x_{i},x_{j})$, $(K_{y})_{i,j}=\kappa_{y}(y_{i},y_{j})$, and $H=I-\frac{1}{N}\mathds{1}\mathds{1}^{T}$ is the centering matrix.